1. 6(x+3)-(x+4)=-11



Found 2 solutions by solver91311, lynnlo:
Answer by solver91311(24713)



Distribute the 6 across the parentheses. Distribute the -1 across the other set of parentheses.

Collect like terms, i.e. 6x - x and 3 - 4.

Add 1 to both sides of the equation.

Multiply both sides by

John

Egw to Beta kai to Sigma
My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism


Answer by lynnlo(4176)

(1)X=-5
(2)K=-6
(3)X=0
(4)X=-4